In an air standard Otto cyclc the maximum and minimum tcmperatures arc 1400
1
Expert's answer
2021-07-26T05:57:02-0400

Compression ratio

"Q_{2-3}=C_v(T_3-T_2)\\
800=0.718(1673-T_2)\\
T_2= 1673-\frac{800}{0.718}=558.8 K\\
\frac{T_2}{T_1}=(\frac{v_1}{v_2})^{\gamma-1}\\
\frac{v_1}{v_2}=(\frac{558.8}{288})^{\frac{1}{0.4}}= 5.243"


Thermal Efficiency

"η= 1-\frac{1}{r_c^{r-1}}=1-\frac{1}{5.243^{0.4}}=48.45 \%"


The ratio of maximum to minimum pressures

"\frac{T_2}{T_1}=(\frac{P_2}{P_1})^{\frac{\gamma-1}{\gamma}}\\
\frac{P_2}{P_1}=(\frac{558.8}{288})^{\frac{1.4}{0.4}}=10.174.............1\\
\frac{P_3}{P_2}=(\frac{T_3}{T_2})=\frac{1673}{558.8}=2.993............2\\
\therefore Eqn \space 1 \space *Eqn \space 2\\
\frac{P_3}{P_1}=10.174*2.993=30.45"